# Amplitude Expansion Phase Field Crystal (APFC) Modeling based Efficient Dislocation Simulations using Fourier Pseudospectral Method

### v2

@@ Abstract
Crystalline defects play a critical role in determining the properties of crystalline solids, underscoring the need for accurate computational methods to study them. Lattice deformation in dislocation simulations, which involves changes in atomic positions, can be described either microscopically by specific atomic configurations or macroscopically by continuum elasticity, each with inherent limitations. The complex amplitude expansion of the phase field crystal (APFC) model provides a mesoscopic approach that bridges these scales. In this paper, we introduce a Fourier pseudospectral method for efficiently solving the APFC model in the context of crystalline defect simulations. This study marks the first application of the Fourier pseudospectral method to the APFC model. The method fully exploits the system’s periodicity and facilitates the implementation of periodic boundary conditions, thanks to its high accuracy and computational efficiency. Numerical experiments conducted on two-dimensional triangular lattices and three-dimensional body-centered cubic lattices for edge dislocation geometry optimization have produced strain field images that align well with by continuum elasticity predictions. The findings demonstrate the potential of the APFC model to accurately capture the complex strain fields associated with dislocations at the mesoscopic scales, a key step toward modeling more intricate crystalline defect structures and dynamics.
keywords:
crystal defects, amplitude expansion, phase field crystal model, Fourier pseudospectral method, elastic strain field
[…]
52B10, 65D18, 68U05, 68U07
@@ 1 Introduction
Crystalline defects, including point defects, dislocations, and cracks, significantly influence the properties of crystalline solids [1, 2, 3, 4, 5]. Accurate and efficient computational methods are crucial for studying these defects, particularly dislocations, which are key to understanding material behavior. While microscopic simulations are often computationally intensive and macroscopic models may lack precision, mesoscopic approaches strike an effective balance by offering coarse-grained degrees of freedom that enhance computational efficiency while maintaining sufficient accuracy, making them especially suitable for dislocation simulations. Prominent mesoscopic models include discrete dislocation dynamics (DDD) [6, 7], coarse-grained molecular dynamics (CGMD) [8, 9], kinetic Monte Carlo (kMC) [10, 11], and phase field crystals (PFC) [12, 13, 14, 15].
@@
The PFC model excels in describing changes in the free energy functional with respect to local atomistic density, providing valuable insights into elasticity and dislocation dynamics. However, its application is limited by the need for fine spatial discretization. To address this, Goldenfeld et al. [16] introduced the complex amplitude expansion of the PFC (APFC) model. This extension, based on the plane wave expansion of the density, allows for the exploration of longer time scales and larger length scales, significantly broadening the model’s applicability in advanced material modeling [17, 18, 19, 20]. The APFC model is particularly useful for studying elastic effects [21, 20], dislocation dynamics [22, 23], and the control of material properties [24, 25]. It can be effectively applied across various atomic descriptions, including those from theoretical modeling [26], atomic simulations [27], and experimental imaging [20].
@@
The primary challenge in APFC models lies in the development and implementation of efficient numerical algorithms, particularly for dislocation simulations, which are crucial in computational materials science. The model is nonlinear with higher-order terms, and the corresponding dynamic equations involve multiple amplitude functions associated with reciprocal lattice vectors. Various numerical methodologies have been proposed to solve the APFC model. The finite difference method was initially employed [16, 28], but it suffered from instability issues, requiring constraints on the time step due to grid spacing limitations [29]. To address this difficulty, semi-implicit methods and spatial adaptivity have been explored. For example, finite element method [30, 24, 31, 32, 33] and Fourier spectral method [34, 35] have been utilized for APFC models. However, finite element methods may not optimally exploit the periodic lattice structures in crystal models. While the Fourier spectral method offers a global representation through Fourier series, it can result in increased free energy if the time step is too large for certain models [36]. Additionally, efficiency remains a concern across these approaches. Therefore, further research is required to comprehensively address these challenges.
The Fourier pseudospectral method [37, 38], closely related to the Fourier spectral method, solves PDEs at collocation points in physical space rather than solving the variational problem. This simplifies the treatment of nonlinear terms and facilitates the evaluation of certain operators. By leveraging fast algorithms like the fast Fourier transform, it significantly accelerates calculations, scaling as $O(N\log N)$, where $N$ represents the number of degrees of freedom. In this paper, we utilize the Fourier pseudospectral method in the spatial domain to fully exploit the periodicity inherent in crystalline structure while enhancing computational efficiency, combined with a semi-implicit discrete scheme for time integration. Our numerical experiments showcase edge dislocations in two-dimensional triangular and three-dimensional body-centered cubic (BCC) lattices. With a careful choice of the initial guess, we solve the APFC model equations using the Fourier pseudospectral method to approximate the amplitude functions. From these solutions, we compute the order parameter and strain field, achieving a distribution consistent with theoretical predictions. Additionally, we analyze the decay of the strain field along a specific direction over a large area, comparing it with standard continuum elasticity theory  [39].
@@ 2 APFC Model
In this section, we provide a derivation of the APFC model from the PFC free energy functional, as well as the corresponding evolution equations for the crystalline lattices considered in this work. The PFC free energy functional  [12, 13, 14, 15] is defined with respect to the dimensionless atomic probability density $n$, which serves as a continuum representation of discrete atomic lattices. The functional is given by,
$F_{n}=\frac{1}{|\Omega|}\int_{\Omega}\big{[}\frac{B_{0}}{2}n^{2}+\frac{B_{x}}{2}n(1+\nabla^{2})^{2}n-\frac{\tau}{3}n^{3}+\frac{v}{4}n^{4}\big{]}\,{\rm d}\,\bm{r},$ (2.1)
where $\Omega$ is the computational domain and $B_{x}\text{,}\;v\text{,}\;\tau$ and $B_{0}$ are parameters that control the phase diagram and properties of underlying system [40]. The amplitude expansion based PFC (APFC) model [41, 42, 16] can be derived by incorporating crystalline information and approximating the density $n$ as a sum of plane waves
$n(\bm{r})=n_{0}(\bm{r})+\sum_{j=1}^{N}\big{[}\eta_{j}(\bm{r})e^{\imath\bm{k_{j}}\cdot\bm{r}}+\eta_{j}^{*}(\bm{r})e^{-\imath\bm{k_{j}}\cdot\bm{r}}\big{]}.$ (2.2)
The set $\{\bm{k_{j}}\}_{j=1}^{N}$ consists of $N$ non-zero reciprocal lattice vectors that represent distinct lattice symmetries, defined as $\bm{k_{j}}=\sum_{i=1}^{N}n_{i}\bm{b_{i}}$, where $\bm{b_{j}}$ are the primitive reciprocal lattice vectors and $n_{i}\in\mathbb{Z}$ are their coefficients. They satisfy the relation $\bm{a_{i}}\cdot\bm{b_{j}}=2\pi\delta_{ij}$, with $\bm{a_{i}}$ being the primitive vectors of the crystal lattice [43]. The variables $\eta_{j}(\bm{r})$ are amplitude functions corresponding to each plane wave, which are crucial for characterizing the state and evolution of the crystal system. Without loss of generality, the average density $n_{0}(\bm{r})$ can be set to zero [20]. We will illustrate two specific examples at the end of this section (cf. Example 2.3).
In the PFC model, the density $n$ satisfies the following dynamic equation, which represents a gradient flow of the energy functional,
$\frac{\partial n}{\partial t}=\nabla^{2}\frac{\delta F_{n}}{\delta n},$ (2.3)
@@
$\frac{\delta F_{n}}{\delta n}=B_{0}n+B_{x}(1+\nabla^{2})^{2}n-\tau n^{2}+vn^{3}.$ (2.4)
@@
QDRG argument
By applying the $\nabla^{2}$ operator to $n$ and using (2.2), we obtain the following result,
$\nabla^{2}n=\sum_{j}e^{\imath\bm{k_{j}}\cdot\bm{r}}(\nabla^{2}+2\imath\bm{k_{j}}\cdot\nabla-|\bm{k_{j}}|^{2})\eta_{j}+\sum_{j}e^{-\imath\bm{k_{j}}\cdot\bm{r}}(\nabla^{2}+2\imath\bm{k_{j}}\cdot\nabla-|\bm{k_{j}}|^{2})\eta^{*}_{j},$ (2.5)
We then utilize the “quick and dirty renormalization group” (QDRG) approach proposed by [42] to capture the essential aspects of the calculations without relying on more rigorous methods. The main idea is to assume that the amplitudes $\eta_{j}(\bm{r})$ remain constant on atomic length scales. As discussed in Remark 2.1, and demonstrated in the corresponding proof in Appendix B.3, we obtain
$\int_{C_{1}}\eta_{j}(\bm{r})e^{\imath\bm{k_{j}}\cdot{\bm{r}}}\,{\rm d}\,\bm{r}\thickapprox\eta_{j}(\bm{r})\int_{C_{1}}e^{\imath\bm{k_{j}}\cdot{\bm{r}}}\,{\rm d}\,\bm{r},$ (2.6)
where $C_{1}$ is the unit cell. Since $e^{\imath\bm{k_{j}}\cdot{\bm{r}}}$ is periodic in the crystal cell, the right hand side of equation (2.6) vanishes unless $\bm{k_{j}}=0$.
Remark 2.1.
To prove (2.6), we consider the integral of its left-hand side over the cell $C_{\varepsilon}$ with $\varepsilon>0$. The following formula holds:
$\Big{\lvert}\frac{1}{\varepsilon^{d}}\int_{C_{\varepsilon}}\eta(\bm{x})e^{\imath\bm{k_{j}}\cdot\frac{\bm{x}}{\varepsilon}}\,{\rm d}\,\bm{x}-\frac{1}{\varepsilon^{d}}\int_{C_{\varepsilon}}I_{\eta}e^{\imath\bm{k_{j}}\cdot\frac{\bm{x}}{\varepsilon}}\,{\rm d}\,\bm{x}\Big{\rvert}=O(\varepsilon),$ (2.7)
@@
$I_{\eta}=\frac{1}{\varepsilon^{d}}\int_{C_{\varepsilon}}\eta(\bm{x})\,{\rm d}\,\bm{x}$ (2.8)
represents the average value of the integral of $\eta(\bm{x})$ over $C_{\varepsilon}$. Here, $\varepsilon^{d}$ denotes the volume of the $C_{\varepsilon}$ cell, and $d$ is the dimension of the model. The complete proof is given in Appendix B.3.
Dynamic equation of amplitude functions
In the APFC framework, the dynamic equation for the density $n$ given by (2.3) is transformed into a dynamic equation for the amplitude functions $\eta_{j}$. This is achieved by multiplying the left-hand side of (2.3) by $e^{-\imath\bm{k}_{j}\cdot\bm{r}}$, integrating over a unit cell $C_{0}$, and normalizing by its volume $|C_{0}|$. By applying the QDRG approach, we can obtain,
$\frac{1}{|C_{0}|}\int_{C_{0}}e^{-\imath\bm{k_{j}}\cdot{\bm{r}}}\cdot\frac{\partial n}{\partial t}\,{\rm d}\,\bm{r}=\frac{1}{|C_{0}|}\int_{C_{0}}\frac{\partial\eta_{j}}{\partial t}+\sum\limits_{\begin{subarray}{c}l=1\\ l\neq j\end{subarray}}^{N}\frac{\partial(\eta_{l}e^{\imath\bm{(k_{l}-k_{j})}\cdot\bm{r}})}{\partial t}+\sum\limits_{j=1}^{N}\frac{\partial(\eta_{j}^{*}e^{-\imath\bm{2k_{j}}\cdot\bm{r}})}{\partial t}\,{\rm d}\,\bm{r}\thickapprox\frac{\partial\eta_{j}}{\partial t}.$ (2.9)
Only the term containing $\eta_{j}$ gives non-zero contribution using the approximation in (2.6), the other terms are zero, since they are multiplied by a periodic function.
Performing the same operation on the right-hand term of (2.3) yields
$\frac{\partial\eta_{j}}{\partial t}=\mathcal{L}_{j}\frac{\delta F_{\eta_{j}}}{\delta\eta_{j}^{*}},$ (2.10)
where $\mathcal{L}_{j}=\nabla^{2}+2\imath\bm{k_{j}}\cdot\nabla-|\bm{k_{j}}|^{2}$.
Remark 2.2.
To simplify the calculation, we employ a long-wavelength limit [44], which allows us to ignore the high-order derivative terms $\nabla^{2}$ and $2\imath\bm{k_{j}}\cdot\nabla$ in $\mathcal{L}_{j}$. As the wave number $\bm{k}\rightarrow 0$, the following result holds
$\lim_{\bm{k}\rightarrow 0}\frac{1}{\varepsilon^{d}}\int_{C_{\varepsilon}}(\nabla^{2}+2\imath\bm{k_{j}}\cdot\nabla)\widehat{\eta_{j}}e^{\imath\bm{k}\cdot\frac{\bm{x}}{\varepsilon}}\,{\rm d}\,\bm{x}=0,$ (2.11)
where $\widehat{\eta_{j}}$ denote the fourier coefficient of $\eta_{j}$. For the simplicity of presentation, we leave the proof of (2.11) in the Appendix B.3.
From the Remark 2.2, we can set
$\mathcal{L}_{j}\thickapprox-|\bm{k_{j}}|^{2}.$ (2.12)
Hence, the dynamic equation of the APFC model (2.10) can be written as
$\frac{\partial\eta_{j}}{\partial t}=-|\bm{k_{j}}|^{2}\frac{\delta F_{\eta_{j}}}{\delta\eta_{j}^{*}},$ (2.13)
@@
$\frac{\delta F_{\eta_{j}}}{\delta\eta_{j}^{*}}=\big{[}B_{0}+B_{x}(\nabla^{2}+2\imath\bm{k_{j}}\cdot\nabla)^{2}+3v(A^{2}-|\eta_{j}|^{2})\big{]}\eta_{j}+\frac{\delta f^{s}}{\delta\eta_{j}^{*}}\big{]}.$ (2.14)
And the free energy functional with respect to $\eta_{j}$ is given by
$F_{\eta_{j}}=\frac{1}{|\Omega|}\int_{\Omega}\Big{[}\frac{B_{0}}{2}A^{2}+\frac{3v}{4}A^{4}+\sum_{j=1}^{N}\big{(}B_{x}|\varsigma_{j}\eta_{j}|^{2}-\frac{3v}{2}|\eta_{j}|^{4}\big{)}+f^{s}({\eta_{j}},{\eta_{j}^{*}})\Big{]}\,\rm{d}r,$ (2.15)
where $\varsigma_{j}:=\nabla^{2}+2ik_{j}\cdot\nabla$. The derivation details of (2.14) and (2.15) are given in the Appendix B.1.
The term $A^{2}$ in (2.14) is defined by
$A^{2}:=2\sum_{j=1}^{N}|\eta_{j}|^{2}.$ (2.16)
It represents a parameter that remains constant within solid and ordered phases but decreases in the presence of defects and interfaces.
The functional $f^{s}(\eta_{j},\eta_{j}^{*})$ corresponds to complex polynomials of $\eta_{j}$ and $\eta_{j}^{*}$, aligned with the appropriate symmetry. In the APFC model (2.13), the crystal lattice symmetry is determined by selecting a suitable set containing $N$ vectors $\{\bm{k_{j}}\}_{j=1}^{N}$. Below, we present two examples of lattice configurations considered in this work.
Example 2.3.
The reciprocal-space vectors used to characterize the crystal symmetry and $f^{s}$ are determined by the chosen lattice symmetry. For simplicity, we present the results for two types of lattices illustrated in Figure 1b, with computational details provided in Appendix B.1.
@@
  • For triangular lattice with $N=3$, the complex polynomial functional in (2.14) reads
    $f^{s}=f^{s}_{\rm{tri}}$ $=-2\tau(\eta_{1}\eta_{2}\eta_{3}+{\rm c.c.}),$
@@
    $\frac{\delta f^{s}_{\rm tri}}{\delta\eta_{j}^{*}}$ =$-2\tau\prod_{i\not=j}^{3}\eta_{i}^{*},\qquad j=1,2,3.$
  • For body-centered cubic (BCC) lattice with $N=6$, the complex polynomial functional in (2.14) reads
    $f^{s}=f^{s}_{\rm bcc}=-$ $2\tau(\eta_{1}^{*}\eta_{2}\eta_{3}+\eta_{2}^{*}\eta_{3}\eta_{5}+\eta_{3}^{*}\eta_{1}\eta_{6}+\eta_{4}^{*}\eta_{5}^{*}\eta_{6}^{*}+{\rm c.c.})$

### v3

@@ Abstract
Crystalline defects critically influence material properties, necessitating accurate simulation methods. Existing approaches, from atomic-scale configurations to continuum elasticity, face inherent limitations in modeling dislocation-induced lattice deformation. The amplitude expansion of the phase field crystal (APFC) model bridges this gap with a mesoscopic description. This paper introduces a computationally efficient Fourier pseudospectral method for solving the APFC equations. The method exploits system periodicity and solution analyticity—the latter’s rigorous proof remaining an open question, as discussed herein—to enable precise implementation of periodic boundary conditions. Numerical experiments on 2D triangular and 3D body-centered cubic lattices demonstrate that the method accurately reproduces the strain fields of edge dislocations, matching continuum theory predictions. These results confirm the APFC model’s potential for capturing complex defect structures at the mesoscale, paving the way for simulating more intricate defect dynamics.
@@ 1 Introduction
Crystalline defects, including point defects, dislocations, and cracks, significantly influence the properties of crystalline solids Lagerlof2018Crystal ; James2017Deformation ; wang2024theoretical ; braun2022higher ; chen2022qm . Accurate and efficient computational methods are crucial for studying these defects, particularly dislocations, which are key to understanding material behavior. While microscopic simulations are often computationally intensive and macroscopic models may lack precision, mesoscopic approaches strike an effective balance by offering coarse-grained degrees of freedom that enhance computational efficiency while maintaining sufficient accuracy, making them especially suitable for dislocation simulations. Prominent mesoscopic models include discrete dislocation dynamics (DDD) lepinoux1987dynamic ; ghoniem1988computer , coarse-grained molecular dynamics (CGMD) rudd1998coarse ; joshi2021review , kinetic Monte Carlo (KMC) lin1999kinetic ; cai2001kinetic , and phase field crystal (PFC) Elder2007Phase-field ; Stefanovic2006Phase-Field .
@@
The PFC model excels in describing changes in the free energy functional with respect to local atomistic density, providing valuable insights into elasticity and dislocation dynamics. However, its application is limited by the need for fine spatial discretization. To address this, Goldenfeld et al. PhysRevE.72.020601 introduced the complex amplitude expansion of the PFC (APFC) model. This extension, based on the plane wave expansion of the density, allows for the exploration of longer time scales and larger length scales, significantly broadening the model’s applicability in advanced material modeling ofori2013complex ; ElderAmplitude2010 ; jreidini2021orientation . The APFC model is particularly useful for studying elastic effects PhysRevE.89.032411 ; ElderAmplitude2010 , dislocation dynamics PhysRevB.97.054113 ; PhysRevLett.121.255501 , and the control of material properties 2017Controlling ; Choudhary01092012 . It can be effectively applied across various atomic descriptions, including those from theoretical modeling de2023amplitude , atomic simulations wang2023pattern , and experimental imaging ElderAmplitude2010 .
@@
The primary challenge in APFC models lies in the development and implementation of efficient numerical algorithms, particularly for dislocation simulations, which are crucial in computational materials science. The model is nonlinear with higher-order terms, and the corresponding dynamic equations involve multiple amplitude functions associated with reciprocal lattice vectors. Various numerical methodologies have been proposed to solve the APFC model. The finite difference method was initially employed PhysRevE.72.020601 ; Yeon07012010 , but it suffered from instability issues, requiring constraints on the time step due to grid spacing limitations provatas2011phase . To address this difficulty, semi-implicit methods and spatial adaptivity have been explored. For example, finite element method 2017Controlling ; 2015Software ; PhysRevE.76.056706 ; PhysRevE.98.033303 and Fourier spectral method PhysRevB.81.214201 ; hirvonen2016multiscale have been utilized for APFC models. However, finite element methods may not optimally exploit the periodic lattice structures in crystal models. While the Fourier spectral method offers a global representation through Fourier series, it can result in increased free energy if the time step is too large for certain models hirvonen2016multiscale . Additionally, efficiency remains a concern across these approaches. Therefore, further research is required to comprehensively address these challenges.
The Fourier pseudospectral method 2011Spectral ; gottlieb1977numerical solves PDEs at collocation points, simplifying the treatment of nonlinear terms. Its computational efficiency, scaling as $\mathcal{O}(N\log N)$ via the fast Fourier transform, is ideal for leveraging the periodicity of crystalline structures. We apply this method spatially, combined with a semi-implicit time integration scheme, to solve the APFC model for edge dislocations in 2D triangular and 3D BCC lattices. With a suitable initial guess, we compute amplitude functions, from which the derived strain fields align with theoretical predictions anderson2017theory . We further validate the results by analyzing the long-range decay of the strain, showing consistency with continuum elasticity. We note that the spectral convergence observed numerically in Section 4 requires a rigorous justification. Such a proof would depend on the analyticity of the exact solution (Assumption 1) and likely the energy dissipation of both the continuous and discrete dynamics. This analysis is beyond the scope of this paper and will be addressed in future work (see Remark 3 and Remark 5).
@@ 2 APFC Model
In this section, we provide a derivation of the APFC model from the PFC free energy functional, as well as the corresponding evolution equations for the crystalline lattices considered in this work. The PFC free energy functional  Elder2007Phase-field ; Stefanovic2006Phase-Field is defined with respect to the dimensionless atomic probability density $n$, which serves as a continuum representation of discrete atomic lattices. The functional is given by
$F_{n}=\frac{1}{\lvert\Omega\rvert}\int_{\Omega}\Big[\frac{B_{0}}{2}n^{2}+\frac{B_{x}}{2}n(1+\nabla^{2})^{2}n-\frac{\gamma}{3}n^{3}+\frac{v}{4}n^{4}\Big]\,{\rm d}\bm{r},$ (1)
where $\Omega$ is the computational domain and $B_{x}\text{,}\;v\text{,}\;\gamma$, and $B_{0}$ are parameters that control the phase diagram and properties of the underlying system Elder2007Phase-field . The APFC model 2006Renormalization ; Athreya2006Renormalization ; PhysRevE.72.020601 can be derived by incorporating crystalline information and approximating the density $n$ as a sum of plane waves
$n(\bm{r},t)=n_{0}(\bm{r},t)+\sum_{j=1}^{M}\big[\eta_{j}(\bm{r},t)e^{\imath\bm{k}_{j}\cdot\bm{r}}+\eta_{j}^{*}(\bm{r},t)e^{-\imath\bm{k}_{j}\cdot\bm{r}}\big].$ (2)
The set $\{\bm{k}_{j}\}_{j=1}^{M}$ consists of $M$ non-zero reciprocal lattice vectors that represent distinct lattice symmetries, defined as $\bm{k}_{j}=\sum_{i=1}^{M}n_{i}\bm{b}_{i}$, where $\bm{b}_{i}$ are the primitive reciprocal lattice vectors and $n_{i}\in\mathbb{Z}$ are their coefficients. They satisfy the relation $\bm{a}_{i}\cdot\bm{b}_{j}=2\pi\delta_{ij}$, with $\bm{a}_{i}$ being the primitive vectors of the crystal lattice zhang2008efficient . The variables $\eta_{j}(\bm{r},t)$ are amplitude functions corresponding to each plane wave, which are crucial for characterizing the state and evolution of the crystal system. Without loss of generality, the average density $n_{0}(\bm{r},t)$ can be set to zero ElderAmplitude2010 . We will illustrate two specific examples at the end of this section (cf. Example 1).
In the PFC model, the density $n$ satisfies the following dynamic equation, which represents a gradient flow of the energy functional
$\frac{\partial n}{\partial t}=\nabla^{2}\frac{\delta F_{n}}{\delta n},$ (3)
@@
$\frac{\delta F_{n}}{\delta n}=B_{0}n+B_{x}(1+\nabla^{2})^{2}n-\gamma n^{2}+vn^{3}.$ (4)
@@
QDRG argument.
Applying the $\nabla^{2}$ operator to $n$ and using (2) yields
$\nabla^{2}n=\sum_{j}e^{\imath\bm{k}_{j}\cdot\bm{r}}(\nabla^{2}+2\imath\bm{k}_{j}\cdot\nabla-\lvert\bm{k}_{j}\rvert^{2})\eta_{j}+\sum_{j}e^{-\imath\bm{k}_{j}\cdot\bm{r}}(\nabla^{2}+2\imath\bm{k}_{j}\cdot\nabla-\lvert\bm{k}_{j}\rvert^{2})\eta^{*}_{j}.$ (5)
We then utilize the “quick and dirty renormalization group” (QDRG) approach proposed by Athreya2006Renormalization to capture the essential aspects of the calculations without relying on more rigorous methods. The main idea is to assume that the amplitudes $\eta_{j}(\bm{r},t)$ remain constant on atomic length scales. As shown in Remark 1, this leads to
$\int_{C_{1}}\eta_{j}(\bm{r},t)e^{\imath\bm{k}_{j}\cdot{\bm{r}}}\,{\rm d}\bm{r}\thickapprox\eta_{j}(\bm{r},t)\int_{C_{1}}e^{\imath\bm{k}_{j}\cdot{\bm{r}}}\,{\rm d}\bm{r},$ (6)
where $C_{1}$ is the unit cell. Since $e^{\imath\bm{k}_{j}\cdot{\bm{r}}}$ is periodic in the crystal cell, the right-hand side of equation (6) vanishes unless $\bm{k}_{j}=0$.
Remark 1
To prove (6), we consider the integral of its left-hand side over the cell $C_{\varepsilon}$ with $\varepsilon>0$. The following formula holds:
$\Big\lvert\frac{1}{\varepsilon^{d}}\int_{C_{\varepsilon}}\eta(\bm{x},t)e^{\imath\bm{k}_{j}\cdot\frac{\bm{x}}{\varepsilon}}\,{\rm d}\bm{x}-\frac{1}{\varepsilon^{d}}\int_{C_{\varepsilon}}I_{\eta}e^{\imath\bm{k}_{j}\cdot\frac{\bm{x}}{\varepsilon}}\,{\rm d}\bm{x}\Big\rvert=O(\varepsilon),$ (7)
@@
$I_{\eta}=\frac{1}{\varepsilon^{d}}\int_{C_{\varepsilon}}\eta(\bm{x},t)\,{\rm d}\bm{x}$ (8)
represents the average value of the integral of $\eta(\bm{x},t)$ over $C_{\varepsilon}$. Here, $\varepsilon^{d}$ denotes the volume of the $C_{\varepsilon}$ cell, and $d$ is the dimension of the model. The complete proof is given in Appendix B.3.
Dynamic equation of amplitude functions.
In the APFC framework, the dynamic equation for the density $n$ given by (3) is transformed into a dynamic equation for the amplitude functions $\eta_{j}$. This is achieved by multiplying the left-hand side of (3) by $e^{-\imath\bm{k}_{j}\cdot\bm{r}}$, integrating over a unit cell $C_{0}$, and normalizing by its volume $\lvert C_{0}\rvert$. By applying the QDRG approach, we can obtain
$\frac{1}{\lvert C_{0}\rvert}\int_{C_{0}}e^{-\imath\bm{k}_{j}\cdot{\bm{r}}}\cdot\frac{\partial n}{\partial t}\,{\rm d}\bm{r}$ $=\frac{1}{\lvert C_{0}\rvert}\int_{C_{0}}\frac{\partial\eta_{j}}{\partial t}+\sum\limits_{\begin{subarray}{c}l=1\\ l\neq j\end{subarray}}^{M}\frac{\partial(\eta_{l}e^{\imath(\bm{k}_{l}-\bm{k}_{j})\cdot\bm{r}})}{\partial t}\,{\rm d}\bm{r}$ (9)
$\quad+\frac{1}{\lvert C_{0}\rvert}\int_{C_{0}}\sum\limits_{j=1}^{M}\frac{\partial(\eta_{j}^{*}e^{-\imath 2\bm{k}_{j}\cdot\bm{r}})}{\partial t}\,{\rm d}\bm{r}\thickapprox\frac{\partial\eta_{j}}{\partial t}.$ (10)
According to the approximation in (6), only the term containing $\eta_{j}$ yields a non-zero contribution because the other terms vanish when multiplied by a periodic function.
Performing the same operation on the right-hand side of (3) leads to
$\frac{\partial\eta_{j}}{\partial t}=\mathcal{L}_{j}\frac{\delta F}{\delta\eta_{j}^{*}},$ (11)
where $\mathcal{L}_{j}=\nabla^{2}+2\imath\bm{k}_{j}\cdot\nabla-\lvert\bm{k}_{j}\rvert^{2}$.
Remark 2
To simplify the calculation, we employ a long-wavelength limit ElderAmplitude2010 , which allows us to ignore the high-order derivative terms $\nabla^{2}$ and $2\imath\bm{k}_{j}\cdot\nabla$ in $\mathcal{L}_{j}$. Specifically, as the wave number $\bm{k}\rightarrow 0$, the following result holds:
$\lim_{\bm{k}\rightarrow 0}\frac{1}{\varepsilon^{d}}\int_{C_{\varepsilon}}(\nabla^{2}+2\imath\bm{k}_{j}\cdot\nabla)\widehat{\eta}_{j}e^{\imath\bm{k}\cdot\frac{\bm{x}}{\varepsilon}}\,{\rm d}\bm{x}=0,$ (12)
where $\widehat{\eta}_{j}$ denote the Fourier coefficient of $\eta_{j}$. For the simplicity of presentation, the proof of (12) is provided in Appendix B.3.
Based on Remark 2, we approximate $\mathcal{L}_{j}\thickapprox-\lvert\bm{k}_{j}\rvert^{2}$. Hence, the dynamic equation of the APFC model (11) becomes
$\frac{\partial\eta_{j}}{\partial t}=-\lvert\bm{k}_{j}\rvert^{2}\frac{\delta F}{\delta\eta_{j}^{*}},$ (13)
@@
$\frac{\delta F}{\delta\eta_{j}^{*}}=\big[B_{0}+B_{x}(\nabla^{2}+2\imath\bm{k}_{j}\cdot\nabla)^{2}+3v(A^{2}-\lvert\eta_{j}\rvert^{2})\big]\eta_{j}+\frac{\delta f^{s}}{\delta\eta_{j}^{*}}.$ (14)
The corresponding free energy functional with respect to $\eta_{j}$ is given by
$F=\frac{1}{\lvert\Omega\rvert}\int_{\Omega}\Big[\frac{B_{0}}{2}A^{2}+\frac{3v}{4}A^{4}+\sum_{j=1}^{M}\big(B_{x}\lvert\varsigma_{j}\eta_{j}\rvert^{2}-\frac{3v}{2}\lvert\eta_{j}\rvert^{4}\big)+f^{s}({\eta_{j}},{\eta_{j}^{*}})\Big]\,{\rm d}\bm{r},$ (15)
where $\varsigma_{j}:=\nabla^{2}+2\imath\bm{k}_{j}\cdot\nabla$. The derivation details of (14) and (15) are given in the Appendix B.1.
The term $A^{2}$ in (14) is defined by
$A^{2}:=2\sum_{j=1}^{M}\lvert\eta_{j}\rvert^{2}.$ (16)
This parameter remains constant within solid and ordered phases but decreases in the presence of defects and interfaces. The functional $f^{s}(\eta_{j},\eta_{j}^{*})$ corresponds to complex polynomials of $\eta_{j}$ and $\eta_{j}^{*}$, aligned with the appropriate symmetry.
Analyticity of solutions.
The regularity of solutions has rarely been addressed in previous studies on phase-field crystal models. In this work, we provide a brief discussion of the analyticity of solutions to (13). At the steady state, the system reduces to a nonlinear elliptic system, for which Morrey 1958Morrey established fundamental results ensuring analyticity of solutions. On the other hand, for the time-dependent case, the Cauchy–Kovalevskaya theorem evans2022partial applies: for parabolic PDE systems with analytic coefficients and analytic, non-characteristic initial data, the solution is analytic over a short time interval. Combining these observations, we obtain the following result.
Theorem 2.1
For the nonlinear parabolic PDE system (13), if the coefficients are analytic and the initial data are analytic and non-characteristic, then the solution $\eta_{j}$ is real analytic in space and time, at least locally in time.
Remark 3
The analyticity established in Theorem 2.1 provides partial justification for applying Fourier pseudospectral methods to the APFC model. To our knowledge, this work presents the first explicit statement of solution analyticity for the APFC system. While Morrey 1958Morrey addressed the steady-state case in elliptic settings, and short-time analyticity for the full parabolic system follows from the Cauchy–Kovalevskaya theorem evans2022partial , long-time analyticity remains an open problem. This persistence may arise from the combination of short-time analyticity and energy dissipation. The numerical experiments in Section 4 further support this regularity property, demonstrating convergence rates consistent with analytic solutions.
Assumption 1
For the nonlinear parabolic PDE system (13), if the coefficients are analytic and the initial data are analytic and non-characteristic, the solutions $\eta_{j}(\bm{x},t)$ remain real analytic in space and time for any $t>0$.
In the APFC model (13), the crystal lattice symmetry is determined by selecting a suitable set of $M$ vectors $\{\bm{k}_{j}\}_{j=1}^{M}$. Below, we present two examples of lattice configurations considered in this work.
Example 1
The reciprocal-space vectors used to characterize the crystal symmetry and $f^{s}$ are determined by the chosen lattice symmetry. For simplicity, we present the results for two types of lattices illustrated in Figure 1, with computational details provided in Appendix B.1.
@@
  • For triangular lattice with $M=3$, the complex polynomial functional $f^{s}$ in (14) reads
    $f^{s}=f^{s}_{\rm{tri}}$ $=-2\gamma(\eta_{1}\eta_{2}\eta_{3}+{\rm c.c.}),$
@@
    $\frac{\delta f^{s}_{\rm tri}}{\delta\eta_{j}^{*}}$ $=-2\gamma\prod_{i\not=j}^{3}\eta_{i}^{*},\qquad j=1,2,3.$
  • For body-centered cubic (BCC) lattice with $M=6$, the complex polynomial functional $f^{s}$ in (14) reads
    $f^{s}=f^{s}_{\rm bcc}=$- $2\gamma(\eta_{1}^{*}\eta_{2}\eta_{3}+\eta_{2}^{*}\eta_{3}\eta_{5}+\eta_{3}^{*}\eta_{1}\eta_{6}+\eta_{4}^{*}\eta_{5}^{*}\eta_{6}^{*}+{\rm c.c.})$